What is the present value of an investment that pays $100 every other year forever when the first cash flow occurs in one year? What is the value if the first cash flow occurs in two years? The opportunity cost of capital is 16% per year.
Answer #1

1.
=100/1.16*1/(1-1/1.16^2)=335.648148148148

2.
=100/1.16^2*1/(1-1/1.16^2)=289.351851851852